OpenStudy (anonymous):

Medal Multiply 2 11/14 and 3 1/15 Express your answer as a mixed number or improper fraction in lowest terms.

OpenStudy (inowalst):

Find common denominators.

OpenStudy (inowalst):

Can you make both denominators the same?

OpenStudy (anonymous):

how would i do that

OpenStudy (inowalst):

My bad. All you have to do is multiply numerator times numerator and denominator times denominator.

OpenStudy (anonymous):

ok 39/14 46/15

OpenStudy (anonymous):

1794/ 210

OpenStudy (anonymous):

897/105

OpenStudy (inowalst):

\[2\frac{ 11 }{ 14 }\times3\frac{ 1 }{ 15 }\] 11 times 1. 14 times 15. Then 2 times 3.

OpenStudy (anonymous):

6 11/210

OpenStudy (inowalst):

Alright, now can you simplify?

OpenStudy (anonymous):

it cant be simplifed

OpenStudy (inowalst):

Okay, then that is your answer.

OpenStudy (anonymous):

do you think it can be reduced?

OpenStudy (inowalst):

\[2\frac{ 11 }{ 15 }\times3\frac{ 1 }{ 15 }=6\frac{ 11 }{ 210 }\] No. 11 is odd. 210 is even. Not possible, unless you want to get a decimal which you dont.

OpenStudy (anonymous):

ok thank you

OpenStudy (inowalst):

You're welcome!